OpenStudy (abhisar):

Genetics Fun! a) A dihybrid plant with incomplete linkage on test cross may produce how many types of plant? b)A test cross of \(\sf F_1\) flies +a/+b produced the following offspring ++/ab = 9% ab/ab = 9% +b/ab =41% a+/ab =41% What will be the distance between linked genes?

OpenStudy (anonymous):

Distance in map units, also called centi-Morgans, is the same as the percent recombination. Based on the genotype of the F1, +a and +b are parental while ++ and ab are recombinant, meaning that there was a break between a and b during crossing over. 9+9 (sum of recombinants)/100 (sum of progeny)= 18/100= 18% or a distance of 18 map units or centi-Morgans.

OpenStudy (abhisar):

That's correct!! But can you tell me the arrangement of genes on the chromosome, cis or trans?

OpenStudy (abhisar):

c:

OpenStudy (anonymous):

They are in repulsion, or trans, because the dominant alleles for the a and b genes are on opposite chromosomes.

OpenStudy (abhisar):

\(\huge \checkmark\)

OpenStudy (anonymous):

:)

OpenStudy (abhisar):

Ok, I am gonna throw one more question c: A dihybrid plant with incomplete linkage on test cross may produce how many types of plant?

OpenStudy (anonymous):

It's the same as with complete dominance, 4 types; only the ratios will be different due to the linkage. The alleles, and therefore the possible phenotypes, do not change.

OpenStudy (abhisar):

That's correct c:

OpenStudy (anonymous):

Please close the questions when they have been answered so others will have the chance to have their questions answered! :)